40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2859     Entry time: Wed Feb 23 17:23:35 2022
Author: Chris 
Type: Electronics 
Category: PSOMA 
Subject: current driver status 

The current driver now in the EE shop for troubleshooting (chassis S1500267/PCB S1500270) looks to be a D1200719-v3 PCB that has been modded to v4. The mod to v4 is also documented in the DCC traveler.

I repeated Zach’s 2017 coarse current adjust knob response test with this unit driving a 25 Ω dummy load, and got slightly different results. There’s a bigger offset at the zero knob setting, and saturation sets in above a knob setting of ~7.

It may be that Zach followed through on his stated plan to mod the circuit with a voltage limiter, like the one added in v7 of the schematic.

Next steps are to check the board for signs of this possible mod, and see if the observed misbehavior can be reproduced.

Other mods in v6/v7 of the circuit address high frequency oscillations and crossover issues, and we may want to implement those as well.
 

Attachment 1: coarseknob.png  67 kB  | Hide | Hide all
coarseknob.png
ELOG V3.1.3-